# High-Capacity and Low-PAPR BICM-OFDM Systems Using Non-Equiprobable and Non-Uniform Constellation Shaping With Clipping and Filtering

### v1

@@ V-B BER Comparison Over AWGN Channel
Figure 12: BER comparison over an AWGN channel for the BICM-OFDM systems without CAF (no PAPR reduction) and with CAF/CNC employing three different PAM constellations.
The BER performance over an AWGN channel for the BICM-OFDM systems with different constellations is compared in Fig. 12, evaluating either without CAF (no PAPR reduction) or with CAF and CNC. As for constellations, Modes A and B are applied to $\bf S1$} and $\bf S2$}, respectively. In addition, the uniform PAM and one-dimensional NUC adopted in ATSC 3.0 are applied to both systems. (Note that the results of S3 are omitted as they exhibit similar performance to S1 over the AWGN channel.) In the case of S1, it can be confirmed that the performance gain achieved by Mode A agrees with that estimated from the BMI values in Fig. 9. The performance is also comparable to that achieved by the one-dimensional NUC of ATSC 3.0.
On the other hand, in the case of S2, Mode B achieves a significant gain, yielding an improvement of approximately $2.35$ dB over the uniform PAM and $0.75$ dB over the NUC when evaluated at a BER of $10^{-5}$. This gain, combined with the performance recovery achieved through CNC, significantly reduces the gap from S1 (without PAPR reduction). It is noteworthy that the performance gap between S1 with the uniform PAM and S2 with Mode B is only $0.50$ dB, while they have a significant difference in terms of PAPR. This suggests that the proposed NENU constellation together with CAF and CNC can significantly reduce PAPR without noticeable SNR penalty over the conventional BICM-OFDM system.
@@ V-C BER Comparison Over Fading Channel
Figure 13: BER comparison over a frequency-selective Rayleigh fading channel for the BICM-OFDM systems without CAF (no PAPR reduction), with CAF/CNC, and with DFT precoding, each employing three different PAM constellations.
Finally, we compare BER performance among the three systems S1, S2, and S3 over a frequency-selective Rayleigh fading channel in Fig. 13. As for constellations, Modes A and C are applied to $\bf S1$} and $\bf S2$}, respectively. In addition, the uniform PAM and one-dimensional NUC adopted in ATSC 3.0 are applied to all three systems for comparison. The performance of S2 is found effective over fading channels, achieving a gain of approximately $2.50$ dB with the proposed NENU constellation. On the other hand, the performance of S3 (employing an MMSE detector) is considerably inferior to that of the other OFDM-based systems. It is known that single-carrier signals experience BER degradation when suboptimal detectors, such as MMSE, are employed at the receiver [49, 50]. The results thus indicate that the performance loss associated with DFT precoding is more substantial than that caused by CAF with CNC. Consequently, for PAPR reduction in OFDM systems utilizing constellation shaping, CAF combined with CNC is superior to DFT precoding in terms of both the PAPR reduction effect and BER performance over fading channels.
@@ VI Conclusion
In this work, we developed high-capacity and low-PAPR OFDM-BICM systems based on one-dimensional NENU constellations for shaping gain along with CAF and CNC for efficient signal peak power reduction without noticeable performance degradation. As a result, the developed systems enjoy significantly lower PAPR while achieving comparable performance with the standard PAM-based BICM-OFDM system. The simulation results have demonstrated that significant performance gain can be achieved over the conventional BICM-OFDM with DFT precoding in terms of both PAPR distribution and BER. In our simulations, we have restricted our attention to the design of $32$-PAM (i.e., $1024$-QAM), but its extension to even higher-order constellations targeting even higher information rates is straightforward, considering the simplicity of the optimization process based on only the two distinct parameters.

### v2

@@ V-B BER Comparison Over AWGN Channel
Figure 12: BER comparison over an AWGN channel for the BICM-OFDM systems without CAF (no PAPR reduction) and with CAF/CNC employing the four different modulation schemes.
Fig. 12 compares the BER performance over an AWGN channel for the BICM-OFDM systems using different modulation schemes, considering both cases without CAF (i.e., no PAPR reduction) and with CAF and CNC. As for the proposed shaping scheme, the constellations $\mathcal{X}_{\sf A}$} and $\mathcal{X}_{\sf B}$} are applied to $\bf S1$} and $\bf S2$}, respectively. Note that the results of S3 are omitted as they exhibit similar performance to S1 over the AWGN channel. In the case of S1, it can be confirmed that the performance gain achieved by the constellation $\mathcal{X}_{\sf A}$} agrees with that estimated from the BMI values in Fig. 9. The resulting BER performance is also comparable to that obtained by the one-dimensional NUC adopted in ATSC 3.0. Although the two-dimensional NUC proposed in[sillekens2022high] achieves the largest shaping gain, it requires much more demapping complexity compared with other modulation schemes based on one-dimensional constellations, as discussed in Section III.
On the other hand, in the case of S2, the constellation $\mathcal{X}_{\sf B}$} achieves a significant gain over the standard uniform PAM. Since $\mathcal{X}_{\sf B}$} is optimized in the presence of CAF/CNC, the BER exhibits a steep decay with increasing SNR, indicating strong robustness to residual clipping noise. As a result, $\mathcal{X}_{\sf B}$} also outperforms both the one-dimensional NUC and the two-dimensional NUC, which are optimized for conventional OFDM systems without PAPR reduction, particularly at low BER levels. These results confirm the effectiveness of optimizing the constellation for the specific OFDM system incorporating nonlinear CAF and CNC operations. The shaping gain achieved by $\mathcal{X}_{\sf B}$}, assisted by the performance recovery through CNC, significantly reduces the gap relative to S1 (without PAPR reduction). It is noteworthy that the performance gap between S1 with the uniform PAM and S2 with $\mathcal{X}_{\sf B}$} is only about $0.50$ dB, while they have a notable difference in terms of PAPR. This suggests that the proposed NENU constellation together with CAF and CNC can significantly reduce PAPR without a noticeable SNR penalty compared to the conventional BICM-OFDM system.
@@ V-C BER Comparison Over Fading Channel
Figure 13: BER comparison over the fading channel for the BICM-OFDM systems without CAF (no PAPR reduction), with CAF/CNC, and with DFT precoding, each employing the four different modulation schemes.
Finally, we compare BER performance among the three systems S1, S2, and S3 over the fading channel (5G NR TDL-B) in Fig. 13. As for the proposed shaping scheme, the constellations $\mathcal{X}_{\sf A}$} is applied to both $\bf S1$} and $\bf S3$}, whereas $\mathcal{X}_{\sf C}$} is applied to $\bf S2$}. It is confirmed that the proposed NENU constellation provides a shaping gain for all systems over the practical fading channel as well. Moreover, the performance of S3 (employing an MMSE detector) is considerably inferior to that of the other OFDM-based systems. It is known that single-carrier signals experience BER degradation when suboptimal detectors, such as MMSE, are employed at the receiver [nasri2009performance, tajer2010diversity]. The results thus indicate that the performance loss associated with DFT precoding is more pronounced than that caused by CAF with CNC. Consequently, for PAPR reduction in OFDM systems utilizing constellation shaping, CAF combined with CNC is more advantageous than DFT precoding from the viewpoint of PAPR as well as BER performance over fading channels.
@@ VI Conclusion
In this work, we developed high-capacity and low-PAPR OFDM-BICM systems based on one-dimensional NENU constellations for shaping gain along with CAF and CNC for efficient signal peak power reduction without noticeable performance degradation. As a result, the developed systems enjoy significantly lower PAPR while achieving comparable performance with the standard PAM-based BICM-OFDM system. The simulation results have demonstrated that significant performance gain can be achieved over the conventional BICM-OFDM with DFT precoding in terms of both PAPR distribution and BER. In our simulations, we have restricted our attention to the design of $32$-PAM (i.e., $1024$-QAM), but its extension to even higher-order constellations targeting even higher information rates is straightforward, considering the simplicity of the optimization process based on only the two shaping parameters.
As final remarks, our constellation design has been optimized for the simulated systems, and its robustness against other channels or system setups should be worth investigating. Furthermore, it is practically important to optimize the overall system by considering the effect of power amplifier nonlinearity in the presence of CAF/CNC, which we leave as future work.